OpenStudy (anonymous):

IF m ANGLE MYZ=2X+5 and m angle XYZ=130 degree find x?

jhonyy9 (jhonyy9):

so if you check it right you will see that mANGLE MYZ mean angle Y and the same mANGLE XYZ will mean the same measure of angle Y - so than this sign that 2x+5=130 2x=125 x=125/2 x=62,5 hope so much that is understandably sure right easy for you good luck bye